OpenStudy (anonymous):

Evaluate: 4 − 49 ÷ 7 + 3^2 −12 2.6 3 6

OpenStudy (anonymous):

4 − 49 ÷ 7 + 3^2

OpenStudy (anonymous):

Sorry

OpenStudy (king):

49/7=7 7+9=16 4-16=-12 -12!

OpenStudy (anonymous):

@King Thanks So Much!